OpenStudy (anonymous):

Determine what the graph of x^2+4xy+y^2-32=0 is. a.) a hyperbola b.) a line c.) an ellipse d.) a parabola

OpenStudy (anonymous):

According to wolframalpha it is a hyperbola. Why?

OpenStudy (anonymous):

Think \[ B^2 - 4 AC \]

OpenStudy (anonymous):

\[ x^2+4xy+y^2-32=0 \\ A x^2 + B xy + Cy^2 -32=0\\ B^2- 4 A C = 4^2 - 4 (1)(1)= 16-4=12>0 \] It is a hyperbola

OpenStudy (anonymous):

Did you understand what I did